The market demand curve for a product is given by

DD: p=200-q

ss:p=50+0.5q


1
Expert's answer
2022-03-01T15:04:01-0500

At equilibrium, "Q_d=Q_s"


"Q_d: =200-q"


"Q_s=50+0.5q"


Hence;

"200-q=50+0.5q"


"150=1.5q"

"q=100"


So, Price is given by;

"200-q"

"=200-100"

"=100"


Therefore, the Price is 100 and the quantity demanded is 100 units